# Mirage: Cross-Embodiment Zero-Shot Policy Transfer with Cross-Painting

### v1

@@ VI-B Physical Experiments
Table III shows the results. We can see that, for both gripper transfer and robot (and gripper) transfer, Mirage achieves strong zero-shot performance, significantly outperforming both baselines. Similar to Tables I and II, the stronger the source policy is, the smaller the performance gap is during transfer. Interestingly, unlike Tables I and II, we do not observe that transferring between robots with the same gripper achieves higher performance than transferring between grippers only. We hypothesize that this is due to the minor scene setup differences that cause a natural drop in policy performance. On the other hand, the failure modes we observe on the different robots or grippers are all very similar to those from the source policy on the source robot, such as not dropping the cup high enough during cup stacking or not keeping the gripper low enough when closing the door of the toaster oven.
Comparing the two baselines, task-specific Diffusion policies achieve stronger performance on the source robot it is trained on, while Octo achieves better success rates on most tasks when the grippers are swapped. Still, Octo tends to miss more grasps or gets stuck, resulting in a lower success rate than Mirage by 20% to 50%. When being evaluated on the UR5, neither baselines are able to achieve any success on the 3 more difficult tasks.
@@ VI-C Sensitivity Analysis of Mirage
Component Factor Variation Success Rate
Visual
Calibration Error
~10 pixels 80%
@@
2 cameras
Source 80%
Target 70%
Background
Different 0%
Control
Gain
@@
x0.5 90%
Offset
8cm 0%
TABLE IV: Sensitivity of Mirage Components on Policy Performance on the Tiger Pick-and-Place Task. For the visual components, we study the effect of camera calibration error, the luminance of inpainted robot rendering, changes in the background between the source and target, and the performance on a 2-camera setup. For the control components, we evaluate sensitivity of Mirage to the control gain of the target robot controller and the $z$-offset in the proprioceptive values.
We examine the sensitivity of policy performance to the various components of Mirage using the Tiger Pick-and-Place task. For the visual components, we study the effect of camera calibration error on the target robot, the luminance of the rendered robot, and background changes between the target and source robots. To study the effect of camera calibration, we add offsets to the masks and inpainted robots. As the calibration error increases, the segmentation masks become less accurate and leave parts of the robot unmasked. This leads to cross-painted images to include remnants of the target robot (see Fig. 3(a)). Table IV shows that large artifacts result in non-negligible policy performance degradation. For the luminance effect, we adjust the inpainted pixels’ luminance by an offset amount, and Table IV shows that the trained policy is relatively robust to it. For the background change, we evaluate Mirage under a different background (black) from what the policy is trained on (light background), and we find that Mirage does not mitigate the effect of background changes. Additionally, we add another camera in the front view (see Fig. 4), train a Diffusion Policy on both cameras’ observations, and apply Mirage to cross-paint both images during test time. Table IV shows that the target embodiment (UR5 + Robotiq) performance is close to the source, indicating that our method is flexible with the number of cameras.
@@
Figure 3: (a) An example of camera calibration error resulting in failure to mask the target robot out; (b) An example of the artifacts introduced due to large changes in camera angles.
Figure 4: The front camera angle in the 2-camera experiment in Table IV. Top: The actual observations of the target robot during the rollout; Bottom: the cross-painted images generated by Mirage.
@@
  • Small Differences: Up to 1 cm in total translation and ${}^{\circ}$ in total rotation mismatches;
  • Medium Differences: Up to 5 cm in total translation and ${}^{\circ}$ in total rotation mismatches;
  • Large Differences: Up to 10 cm in total translation and ${}^{\circ}$ in total rotation mismatches;
@@
Small (1 cm + ${}^{\circ}$) 90% 90% 90% 90%
Medium (5 cm + ${}^{\circ}$) 60% 80% 50% 80%
Large (10 cm + ${}^{\circ}$) 30% 50% 20% 40%
TABLE V: Mirage Effectiveness of Camera Reprojection. We evaluate Mirage with and without the camera reprojection on the Tiger Pick-and-Place task with 3 levels of camera pose changes: $\leq$ 1 cm and $\leq$ ${}^{\circ}$, $\leq$ 5 cm and $\leq$ ${}^{\circ}$, $\leq$ 10 cm and $\leq$ ${}^{\circ}$. Results suggest that, when there are medium to large differences between camera poses, applying camera reprojection is beneficial.
@@
Table V shows that applying camera reprojection has a 20-30% improvement in the success rates when there are medium to large differences between the camera poses of the source and the target. When the difference increases, there are fewer points projected back to the source image frame, resulting in more missing pixels being inpainted. This leads to blurriness and artifacts close to the boundary of the images (see Fig. 3(b)), making it more likely for the policy to miss the grasp.
@@ Acknowledgments
This research was performed at the AUTOLab at UC Berkeley in affiliation with the Berkeley AI Research (BAIR) Lab, and the CITRIS “People and Robots” (CPAR) Initiative, and in collaboration with Google DeepMind. The authors are supported in part by donations from Google, Toyota Research Institute, and equipment grants from NVIDIA. L.Y. Chen is supported by the National Science Foundation (NSF) Graduate Research Fellowship Program under Grant No. 2146752. We thank Zehan Ma, Simeon Adebola, Max Fu, Ethan Qiu, and Roy Lin for helping with physical experiments, Karl Pertsch and Sudeep Dasari for support with Octo model finetuning, and Will Panitch and Qianqian Wang for valuable feedback.
@@
  • Section C: Additional experiments using diffusion models to improve the visual qualities of the cross-painted robot.
Appendix A Delta Control and Trajectory Playback
In Section IV-B Table I, we see that target robots can achieve high task success rates by querying the source robot policy and referencing the source robot for desired poses. Specifically, we use a blocking controller with absolute pose commands. In this section, we compare and experiment with executing the same delta action from the policy directly on the target robot with a non-blocking controller. Table VI shows that there is a significant drop in performance, indicating that the difference in the forward dynamics between robots cannot be ignored when transferring policies and that leveraging a blocking controller on the target robot is an effective way to mitigate this difference.
Additionally, we compare querying a source robot policy with playing back successful source robot trajectories. Since we have shown that executing the same delta actions as the source robot does not transfer well, we use the achieved poses by the source robots in the demonstrated trajectories paired with blocking Cartesian controllers on the target robots. Table VII shows the results. We see that, compared with Table I, the performances are similar to those using the source robot policies. This suggests that these tasks can be achieved by those target robots using the same strategy. In fact, this is similar to the data generation strategy adopted by MimicGen [69], in which the authors roll out the (transformed) trajectories on the target robots and filter out those unsuccessful ones and then train a model on the rest. Comparing playing back human demonstration trajectories and querying policies on the source robot, we observe that in some cases where the learned policies are brittle or suboptimal, they leave less room for error and are less robust to changes in the gripper or tracking errors. On the other hand, good learned policies can potentially adapt to tracking errors, adjust their poses, and retry grasps if necessary, thus outperforming trajectory playback in some cases.
@@ Appendix A Delta Control and Trajectory Playback
TABLE VI: State-Based Transfer Experiment with Delta Actions. We evaluate state-based policies trained on a Franka robot for each task across five different robots equipped either with the original Franka gripper or with each target robot’s default gripper. The target robot executes the same delta Cartesian action as the source Franka robot instead of reaching the same absolute pose.
@@
TABLE VII: Success Rates of Target Robots Playing Back the Source Robot Trajectories. Instead of executing the source robot policy, we evaluate the success rates of target robots performing trajectory playback. Specifically, we use the robot’s blocking controller with the achieved pose of the source robot as the target pose instead of performing the same delta actions.
@@ Appendix B Additional Implementation Details of Mirage
In this section, we detail the inpainting step in the cross-painting process of Mirage, and its sensitivity to calibration errors. Once we have used the segmentation mask provided by Gazebo to black out the target robot, we use Fast Marching [100] algorithm to fill in the black pixels. Specifically, we use a neighborhood radius of 3 pixels and the Telea inpainting method, which is based on a fast marching algorithm [100]. While the filled-in pixels appear blurry, they generally match the color scheme of the rest of the image when the calibration error is low. However, with higher calibration errors, the segmentation masks become less accurate, resulting in remnants of the target robot in the cross-painted images (see Fig. 3(a)). To mitigate this issue, we dilate the segmentation masks to ensure that the target robot is completely blacked out even with calibration errors. With moderate to low calibration errors, we dilate the robot arm segmentation mask with a 3x3 kernel for 20 iterations and the gripper segmentation mask for 10 iterations. However, when the calibration error is higher, we dilate the arm and gripper segmentation masks by 40 and 20 iterations respectively. The gripper mask is dilated significantly less than the robot arm mask because the gripper is located near the objects we perform tasks on. Too much dilation would blur the objects that the robot needs to interact with, leading to worse performance.
Appendix C Improving Visual Qualities with Diffusion Models
Figure 5: Robot cross-painting with diffusion generation. We utilize a ControlNet to transfer the pre-trained Stable Diffusion into the robotics domain. The upper part presents the training phase, in which we finetune the parameters of ControlNet while keeping the Stable Diffusion frozen using a masked reconstruction objective. The bottom part presents the inference phase, where a URDF.
Considering the advanced state of diffusion models in generating high-quality images, it is natural to ask whether we can leverage diffusion models to bridge the visual gaps. Specifically, there is still a small visual gap between the rendered URDF $o^{\mathcal{T}\rightarrow\mathcal{S}}$ and the oracle image $o^{\mathcal{S}}$. However, since the large-scale image-text pair LAION dataset does not contain many robotics images, pre-trained Stable Diffusion [81] models suffer from generalizing to robotics domain zero-shot.
We explore using diffusion models to improve the visual quality of the URDF cross-painted images of Mirage. Specifically, instead of directly finetuning the original Stable Diffusion model, we tune a ControlNet [118] to bridge this gap, as shown in Fig. 5. Additionally, we adopt a self-supervised masking-reconstruction training paradigm [37]. Specifically, during training, the ControlNet takes in a masked image of the source robot $o^{\mathcal{S}}$ and the goal is to reconstruct $o^{\mathcal{S}}$. During inference, we condition the ControlNet on a URDF cross-painted image $o^{\mathcal{T}\rightarrow\mathcal{S}}$ but without the fast marching inpaint step (i.e., leaving the masked out pixels that are not covered by the source robot as black) to generate a visually improved image that is ideally closer to $o^{\mathcal{S}}$.
We qualitatively experiment with and without adding the diffusion model on top of the URDF cross-painted images. While we find that the robot looks visually more realistic than the analytically rendered images, the policy performance is about the same. Since diffusion generation takes a significant amount of time at each time step, we opt not to integrate that part into Mirage pipeline. We leave further investigation of the potential benefit of diffusion models to future work.

### v3

@@ VI-B Physical Experiments
Additionally, similar to Tables I and II, the stronger the source policy is, the smaller the performance gap is during transfer. Interestingly, unlike Tables I and II, we do not observe that transferring between robots with the same gripper achieves higher performance than transferring between grippers only. We hypothesize that this is due to the minor scene setup differences that cause a natural drop in policy performance. On the other hand, the failure modes we observe on the different robots or grippers are all very similar to those from the source policy on the source robot, such as not dropping the cup high enough during cup stacking or not keeping the gripper low enough when closing the door of the toaster oven.
We also evaluate Mirage on first-person images. We select the “Tiger” and “Drawer” tasks and train a diffusion policy with wrist camera inputs and a diffusion policy that take both first-person and third-person camera inputs. Both models use proprioceptive inputs. For the latter, Mirage applies cross-painting to both images. From Table V, we can see that only Mirage is able to achieve high success rates on the target embodiments. We hypothesize that this is because there is a large visual difference between the 2 grippers (see Figure 3), which confuses the policy if cross-painting is not applied to bridge the gap. This shows that Mirage is flexible with the number of cameras as well as the camera view.
@@ VI-C Sensitivity Analysis of Mirage
{tblr}
row2 = c, row3 = c, row4 = c, row5 = c, row6 = c, row7 = c, row8 = c, row9 = c, row10 = c, row11 = c, cell11 = c, cell12 = c, cell13 = c, cell21 = r=4, cell22 = r=2, cell42 = r=2, cell61 = r=4, cell62 = r=2, cell82 = r=2, vline2-4 = 1-13, vline4 = 3-12, vline3-4 = 4,6,8,10,13, hline1-2,6,10 = -, hline4,8 = 2-4, Component & Factor Variation Success Rate
Visual Calibration Error ~10 pixels 80%
@@
Control Gain x0.5 90%
@@
Offset 2 cm 70%
5 cm 20%
TABLE VI: Sensitivity of Mirage Components on Policy Performance on the Tiger Pick-and-Place Task. For the visual components, we study the effect of camera calibration error and the luminance of inpainted robot rendering. For the control components, we evaluate sensitivity of Mirage to the control gain of the target robot controller and the $z$-offset in the proprioceptive values.
We examine the sensitivity of policy performance to the various components of Mirage using the Tiger-in-Bowl task. For the visual components, we study the effect of camera calibration error on the target robot, the quality of the robot rendering, and background changes between the target and source robots. To study the effect of camera calibration and URDF inaccuracy, we add offsets to the masks and inpainted robots. As the calibration error increases, the segmentation masks become less accurate and leave parts of the robot unmasked. This leads to cross-painted images that include remnants of the target robot (see Fig. 4(a)). Table VI shows that large artifacts result in non-negligible policy performance degradation. For the sensitivity of the renderer quality, we adjust the inpainted pixels’ luminance by an offset amount, and Table VI shows that the trained policy is relatively robust to it.
@@
Figure 4: (a) An example of camera calibration error resulting in failure to mask all of the target robot out; (b) An example of the artifacts introduced due to large changes in camera angles.
@@
  • Small Differences: Up to 1 cm in total translation and 5 in total rotation mismatches;
  • Medium Differences: Up to 5 cm in total translation and 15 in total rotation mismatches;
  • Large Differences: Up to 10 cm in total translation and 15 in total rotation mismatches;
@@
Small (1 cm + 5) 90% 90% 90% 90%
Medium (5 cm + 15) 60% 80% 50% 80%
Large (10 cm + 15) 30% 50% 20% 40%
TABLE VII: Mirage Effectiveness of Camera Reprojection. We evaluate Mirage with and without the camera reprojection on the Tiger Pick-and-Place task with 3 levels of camera pose changes: $\leq$ 1 cm and $\leq$ 5, $\leq$ 5 cm and $\leq$ 15, $\leq$ 10 cm and $\leq$ 15. Results suggest that, when there are medium to large differences between camera poses, applying camera reprojection is beneficial.
@@
Table VII shows that applying camera reprojection has a 20-30% improvement in the success rates when there are medium to large differences between the camera poses of the source and the target. When the difference increases, there are fewer points projected back to the source image frame, resulting in more missing pixels being inpainted. This leads to blurriness and artifacts close to the boundary of the images (see Fig. 4(b)), making it more likely for the policy to miss the grasp.
@@ Acknowledgments
This research was performed at the AUTOLab at UC Berkeley in affiliation with the Berkeley AI Research (BAIR) Lab, and the CITRIS “People and Robots” (CPAR) Initiative, and in collaboration with Google DeepMind. The authors are supported in part by donations from Google, Toyota Research Institute, and equipment grants from NVIDIA. L.Y. Chen is supported by the National Science Foundation (NSF) Graduate Research Fellowship Program under Grant No. 2146752. We thank Zehan Ma, Simeon Adebola, Max Fu, Ethan Qiu, and Roy Lin for helping with physical experiments, Karl Pertsch and Sudeep Dasari for support with Octo model fine-tuning, Ajay Mandlekar for support with Robosuite, and Will Panitch and Qianqian Wang for valuable feedback.
@@
TABLE VIII: State-Based Transfer Experiment with Delta Actions. We evaluate state-based policies trained on a Franka robot for each task across five different robots equipped either with the original Franka gripper or with each target robot’s default gripper. The target robot executes the same delta Cartesian action as the source Franka robot instead of reaching the same absolute pose.
@@
TABLE IX: Success Rates of Target Robots Playing Back the Source Robot Trajectories. Instead of executing the source robot policy, we evaluate the success rates of target robots performing trajectory playback. Specifically, we use the robot’s blocking controller with the achieved pose of the source robot as the target pose instead of performing the same delta actions.
Appendix A Delta Control and Trajectory Playback
In Section IV-B Table I, we see that target robots can achieve high task success rates by querying the source robot policy and referencing the source robot for desired poses. Specifically, we use a blocking controller with absolute pose commands. In this section, we compare and experiment with executing the same delta action from the policy directly on the target robot with a non-blocking controller. Table VIII shows that there is a significant drop in performance, indicating that the difference in the forward dynamics between robots cannot be ignored when transferring policies and that leveraging a blocking controller on the target robot is an effective way to mitigate this difference.
Additionally, we compare querying a source robot policy with playing back successful source robot trajectories. Since we have shown that executing the same delta actions as the source robot does not transfer well, we use the achieved poses by the source robots in the demonstrated trajectories paired with blocking Cartesian controllers on the target robots. Table IX shows the results. We see that, compared with Table I, the performances are similar to those using the source robot policies. This suggests that these tasks can be achieved by those target robots using the same strategy. In fact, this is similar to the data generation strategy adopted by MimicGen [68], in which the authors roll out the (transformed) trajectories on the target robots and filter out those unsuccessful ones and then train a model on the rest. Comparing playing back human demonstration trajectories and querying policies on the source robot, we observe that in some cases where the learned policies are brittle or suboptimal, they leave less room for error and are less robust to changes in the gripper or tracking errors. On the other hand, good learned policies can potentially adapt to tracking errors, adjust their poses, and retry grasps if necessary, thus outperforming trajectory playback in some cases.
@@ Appendix B Additional Implementation Details of Mirage
In this section, we detail the inpainting step in the cross-painting process of Mirage, and its sensitivity to calibration errors. Once we have used the segmentation mask provided by Gazebo to black out the target robot, we use Fast Marching [98] algorithm to fill in the black pixels. Specifically, we use a neighborhood radius of 3 pixels and the Telea inpainting method, which is based on a fast marching algorithm [98]. While the filled-in pixels appear blurry, they generally match the color scheme of the rest of the image when the calibration error is low. However, with higher calibration errors, the segmentation masks become less accurate, resulting in remnants of the target robot in the cross-painted images (see Fig. 4(a)). To mitigate this issue, we dilate the segmentation masks to ensure that the target robot is completely blacked out even with calibration errors. With moderate to low calibration errors, we dilate the robot arm segmentation mask with a 3x3 kernel for 20 iterations and the gripper segmentation mask for 10 iterations. However, when the calibration error is higher, we dilate the arm and gripper segmentation masks by 40 and 20 iterations respectively. The gripper mask is dilated significantly less than the robot arm mask because too much dilation would blur the objects that the robot needs to interact with, leading to worse performance.